# An Exact Solution Algorithm for the Bi-Level Optimization Problem of Electric Vehicles Charging Station Placement

### v1

@@ 4.2.1 Outer Approximation for Value Function Cuts
We will derive the OA around $(\mathbf{y}^{1},\mathbf{v}^{1},\mathbf{x}^{1})$. Note that $L$ is composed of three parts, and only the first part is nonlinear. We will derive the OA just for the nonlinear part since the OA of a linear term is equal to itself. If we consider just the left-hand side of constraint (9j) and derive the OA for that:
$OA\left(\sum_{(i,j)\in A}\int_{0}^{x_{ij}}t_{ij}(x)\,dx\right)\leq\sum_{(i,j)\in A}\int_{0}^{x_{ij}}t_{ij}(x)\,dx$ (11)
$\sum_{(i,j)\in A}\int_{0}^{x_{ij}^{1}}t_{ij}(x)\,dx+\sum_{(i,j)\in A}t_{ij}(x^{1})\cdot(x-x^{1})\leq\sum_{(i,j)\in A}\int_{0}^{x_{ij}}t_{ij}(x)\,dx$ (12)
Equation (12) is separable by link $(i,j)$. Let’s $\eta_{ij}$ the link specific OA of equation (12), we can write:
$\int_{0}^{x_{ij}^{1}}t_{ij}(x)\,dx+t_{ij}(x^{1})\cdot(x_{ij}-x_{ij}^{1})\leq\eta_{ij}\quad\forall x^{1}$ (13)
@@
Both sides of constraint (9j) are nonlinear, but the right-hand side is a constant since it is the answer of the objective function with UE optimal solution, $\mathbf{x}^{\mathrm{f}},\mathbf{v}^{\mathrm{f}}$, so we will derive the OA just for the left-hand side term and just the nonlinear component.
$OA\left(\sum_{(i,j)\in\mathcal{A}}\int_{0}^{x_{ij}}t_{ij}(x)\,dx\right)+\sum_{\pi\in\Pi(\mathbf{y})}\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}}t\,dv\\[6.0pt] -\left(\sum_{(i,j)\in\mathcal{A}}\int_{0}^{x_{ij}^{f}}t_{ij}(x)\,dx+\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}^{\mathrm{f}}}t\,dv\right)-\sum_{l\in\mathcal{C}}M_{l}(1-y_{l})\leq 0$ (14)
If we combine equation (13) and (14):
$\sum_{(i,j)\in A}\eta_{ij}+\sum_{\pi\in\Pi(\mathbf{y})}\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}}t\,dv\leq\sum_{(i,j)\in A}\int_{0}^{x_{ij}^{\mathrm{f}}}t_{ij}(x)\,dx+\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}^{\mathrm{f}}}t\,dv+\sum_{l\in\mathcal{C}}M_{l}(1-y_{l})$ (15)
Therefore, we have approximated cut (9j) in problem (9) with the combination of cuts (13) and (15). The final linearized HPR will be the following:
$\min_{\mathbf{y,x,v},\boldsymbol{\eta}}\quad$ $-\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}}p\,dv+w\cdot\sum_{(r,s)}\left(D_{rs}-\sum_{\pi\in\Pi_{rs}(\mathbf{y})}h_{\pi}\right)$ (16a)
@@
$y_{l}\in\{0,1\},$ $\quad\forall l\in\mathcal{C}$ (16c)
@@
$v_{l}=\sum_{\pi\in\Pi(\mathbf{y})}\delta_{l}^{\pi}h_{\pi},$ $\quad\forall l\in\mathcal{C}$ (16g)
@@
$\sum_{ij\in\mathcal{A}}\eta_{ij}+\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}}t\,dv$
$\leq\sum_{ij\in\mathcal{A}}\int_{0}^{x_{ij}^{\mathrm{f}}}t_{ij}(x)\,dx+\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}^{\mathrm{f}}}t\,dv+\sum_{l\in\mathcal{C}}M_{l}(1-y_{l})$ $\quad\forall\,(\mathbf{x}^{\mathrm{f}},\mathbf{v}^{\mathrm{f}})\in\mathcal{F}$ (16k)
@@ 5.2 Column Generation
Formulation (16) contains one path flow variable $h_{\pi}$ per feasible EV path $\pi\in\Pi(\mathbf{y})$. In realistic networks, $|\Pi(\mathbf{y})|$ is a large set. Following the approach used by [ReyLevin2024], we apply the CG method to efficiently solve the linear programming (LP) relaxation. We consider a restricted set of feasible paths $\bar{\Pi}\subseteq\Pi(\mathbf{y})$, and denote the corresponding restricted path sets as $\bar{\Pi}_{rs}\subseteq\Pi_{rs}(\mathbf{y})$ for all OD pairs $(r,s)\in\mathcal{Z}^{2}$. Problem (16) can be written as follows:
$\min_{\mathbf{y},\mathbf{x},\mathbf{v},\mathbf{h},\boldsymbol{\eta}}\quad$ $-\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}}p\,dv+w\cdot\sum_{(r,s)}\left(D_{rs}-\sum_{\pi\in\bar{\Pi}_{rs}}h_{\pi}\right)$ (19a)
@@
$v_{l}\leq M\cdot y_{l},$ $\quad\forall l\in\mathcal{C}$ (19c)
$x_{ij}=\sum_{\pi\in\bar{\Pi}}\delta_{ij}^{\pi}h_{\pi},$ $\quad\forall(i,j)\in\mathcal{A}$ (19d)
$\sum_{\pi\in\bar{\Pi}_{rs}}h_{\pi}=d_{rs}(\mathbf{y}),$ $\quad\forall(r,s)\in\mathcal{Z}^{2}$ (19e)
$v_{l}=\sum_{\pi\in\bar{\Pi}}\delta_{l}^{\pi}h_{\pi},$ $\quad\forall l\in\mathcal{C}$ (19f)
$\int_{0}^{x_{ij}^{1}}t_{ij}(x)\,dx+t_{ij}(x^{1})(x_{ij}-x_{ij}^{1})\leq\eta_{ij},$ $\quad\forall x^{1}$ (19g)
$\sum_{(i,j)\in\mathcal{A}}\eta_{ij}+\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}}t\,dv$
$\leq\sum_{(i,j)\in\mathcal{A}}\int_{0}^{x_{ij}^{\mathrm{f}}}t_{ij}(x)\,dx+\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}^{\mathrm{f}}}t\,dv+\sum_{l\in\mathcal{C}}M_{l}(1-y_{l}),$ $\quad\forall(\mathbf{x}^{\mathrm{f}},\mathbf{v}^{\mathrm{f}})$ (19h)
$y_{l}\in[0,1],\quad h_{\pi}\geq 0,$ $\quad\forall l\in\mathcal{C},\;\pi\in\bar{\Pi}$ (19i)
@@
Since the objective function aims to maximize path flows $h_{\pi}$ and charging flows $v_{l}$, we reformulate the corresponding constraints as inequalities to expose the sign of their associated dual variables. Specifically, the demand satisfaction constraint (19e) becomes $\sum_{\pi\in\bar{\Pi}_{rs}}h_{\pi}\leq d_{rs}(\mathbf{y})$, the link flow definition (19d) becomes $\sum_{\pi\in\bar{\Pi}}\delta_{ij}^{\pi}h_{\pi}\leq x_{ij}$, and the charging flow definition (19f) becomes $\sum_{\pi\in\bar{\Pi}}\delta_{l}^{\pi}h_{\pi}\geq v_{l}$. We derive dual variables associated with the constraints to determine the reduced cost of adding new paths to the LP. Denote by $\sigma_{rs}\geq 0$ the dual variables for the demand satisfaction constraints (19e), and by $\zeta_{ij}\geq 0$ the dual variables associated with link flow constraints (19d). Given an OD pair $(r,s)\in\mathcal{Z}^{2}$ and a path $\pi\in\Pi_{rs}^{\mathrm{f}}$, the reduced cost $c_{\pi}$ for the path flow variable $h_{\pi}$ is computed as follows:
$c_{\pi}=-\sigma_{rs}+\sum_{(i,j)\in\mathcal{A}}\delta_{ij}^{\pi}\zeta_{ij}-\sum_{l\in\mathcal{C}}\delta_{l}^{\pi}\gamma_{l},$ (20)
@@
where $\gamma_{l}\leq 0$ represents the dual variables for the charging station flow constraints (19f). Thus, to identify new paths that have negative reduced cost, one solves, for each OD pair $(r,s)$, a battery constrained shortest path problem with link costs given by dual variables $\zeta_{ij}$ and charging costs $\gamma_{l}$. Since we have the expanded network we can solve the constrained shortest path problem (CSPP) efficiently. If paths with negative reduced costs are found, these are introduced into the restricted path set $\bar{\Pi}$ to refine the problem iteratively. This process continues until no paths with negative reduced cost are found. Algorithm 2 shows the CG steps to iteratively solve problem (19) using a limited set of paths and updates it by identifying and adding new paths with negative reduced cost.

### v2

@@ 4.2.1 Outer Approximation for Value Function Cuts
We will derive the OA around $(\mathbf{y}^{1},\mathbf{v}^{1},\mathbf{x}^{1})$. Note that $L$ is composed of three parts: the first two parts are nonlinear, while the third part (charging cost) is linear. We will derive the OA for both nonlinear parts since the OA of a linear term is equal to itself. For the first term, if we consider the left-hand side of constraint (9j) and derive the OA:
$OA\left(\sum_{(i,j)\in\mathcal{A}}\int_{0}^{x_{ij}}t_{ij}(x)\,dx\right)\leq\sum_{(i,j)\in\mathcal{A}}\int_{0}^{x_{ij}}t_{ij}(x)\,dx$ (11)
The OA of a convex function $f(x)$ at a point $x^{1}$ is given by the first-order Taylor expansion: $f(x^{1})+\nabla f(x^{1})^{\top}(x-x^{1})\leq f(x)$. Applying this to the link travel time integral, where the derivative of $\int_{0}^{x_{ij}}t_{ij}(x)\,dx$ with respect to $x_{ij}$ is $t_{ij}(x_{ij})$, we obtain:
$\sum_{(i,j)\in\mathcal{A}}\int_{0}^{x_{ij}^{1}}t_{ij}(x)\,dx+\sum_{(i,j)\in\mathcal{A}}t_{ij}(x_{ij}^{1})\cdot(x_{ij}-x_{ij}^{1})\leq\sum_{(i,j)\in\mathcal{A}}\int_{0}^{x_{ij}}t_{ij}(x)\,dx$ (12)
Equation (12) is separable by link $(i,j)$, meaning the inequality can be decomposed into independent terms for each link. To facilitate this decomposition, we introduce auxiliary variables $\eta_{ij}$ for each link $(i,j)\in\mathcal{A}$, which serve as upper bounds on the true integral $\int_{0}^{x_{ij}}t_{ij}(x)\,dx$. By requiring that the OA at any linearization point $x_{ij}^{1}$ is bounded by $\eta_{ij}$, we can write the link-specific OA constraints as:
$\int_{0}^{x_{ij}^{1}}t_{ij}(x)\,dx+t_{ij}(x_{ij}^{1})\cdot(x_{ij}-x_{ij}^{1})\leq\eta_{ij}\quad\forall x_{ij}^{1}$ (13)
These constraints ensure that $\eta_{ij}\geq\int_{0}^{x_{ij}}t_{ij}(x)\,dx$ for all $(i,j)\in\mathcal{A}$, and consequently:
$\sum_{(i,j)\in\mathcal{A}}\eta_{ij}\geq OA\left(\sum_{(i,j)\in\mathcal{A}}\int_{0}^{x_{ij}}t_{ij}(x)\,dx\right)$ (14)
Similarly, for the station congestion term, we derive the OA:
$OA\left(\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}}t_{l}(v)\,dv\right)\leq\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}}t_{l}(v)\,dv$ (15)
Applying the first-order Taylor expansion, where the derivative of $\int_{0}^{v_{l}}t_{l}(v)\,dv$ with respect to $v_{l}$ is $t_{l}(v_{l})$:
$\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}^{1}}t_{l}(v)\,dv+\sum_{l\in\mathcal{C}}t_{l}(v_{l}^{1})\cdot(v_{l}-v_{l}^{1})\leq\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}}t_{l}(v)\,dv$ (16)
Equation (16) is separable by station $l$. We introduce auxiliary variables $\psi_{l}$ for each station $l\in\mathcal{C}$, which serve as upper bounds on the true integral $\int_{0}^{v_{l}}t_{l}(v)\,dv$. By requiring that the OA at any linearization point $v_{l}^{1}$ is bounded by $\psi_{l}$, we can write the station-specific OA constraints as:
$\int_{0}^{v_{l}^{1}}t_{l}(v)\,dv+t_{l}(v_{l}^{1})\cdot(v_{l}-v_{l}^{1})\leq\psi_{l}\quad\forall v_{l}^{1}$ (17)
These constraints ensure that $\psi_{l}\geq\int_{0}^{v_{l}}t_{l}(v)\,dv$ for all $l\in\mathcal{C}$, and consequently:
$\sum_{l\in\mathcal{C}}\psi_{l}\geq OA\left(\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}}t_{l}(v)\,dv\right)$ (18)
We now show how to combine these results to linearize constraint (9j). Recall that the value function cut requires:
$L(\mathbf{y},\mathbf{v}(\mathbf{h}),\mathbf{x}(\mathbf{h}))\leq L(\mathbf{y},\mathbf{v}^{\mathrm{f}},\mathbf{x}^{\mathrm{f}})+\sum_{l\in\mathcal{C}}M_{l}(1-y_{l})$ (19)
@@
Both sides of constraint (9j) are nonlinear, but the right-hand side is a constant since it is evaluated at the UE optimal solution $(\mathbf{x}^{\mathrm{f}},\mathbf{v}^{\mathrm{f}})$. Therefore, we derive the OA only for the left-hand side, and only for the nonlinear components (the charging cost term $\sum_{\pi\in\Pi(\mathbf{y})}c^{\pi}h_{\pi}$ is already linear). Expanding the left-hand side using the definition of $L(\cdot)$ and applying OA to the nonlinear terms:
$OA\left(\sum_{(i,j)\in\mathcal{A}}\int_{0}^{x_{ij}}t_{ij}(x)\,dx\right)+OA\left(\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}}t_{l}(v)\,dv\right)+\sum_{\pi\in\Pi(\mathbf{y})}c^{\pi}h_{\pi}\\[6.0pt] \leq\sum_{(i,j)\in\mathcal{A}}\int_{0}^{x_{ij}^{\mathrm{f}}}t_{ij}(x)\,dx+\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}^{\mathrm{f}}}t_{l}(v)\,dv+\sum_{\pi\in\Pi(\mathbf{y})}c^{\pi}h_{\pi}^{\mathrm{f}}+\sum_{l\in\mathcal{C}}M_{l}(1-y_{l})$ (20)
From inequalities (14) and (18), we have established that the auxiliary variables provide upper bounds on the OA terms:
$OA\left(\sum_{(i,j)\in\mathcal{A}}\int_{0}^{x_{ij}}t_{ij}(x)\,dx\right)\leq\sum_{(i,j)\in\mathcal{A}}\eta_{ij},\quad OA\left(\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}}t_{l}(v)\,dv\right)\leq\sum_{l\in\mathcal{C}}\psi_{l}$ (21)
Substituting these bounds into (20), we obtain the linearized value function cut:
$\sum_{(i,j)\in\mathcal{A}}\eta_{ij}+\sum_{l\in\mathcal{C}}\psi_{l}+\sum_{\pi\in\Pi(\mathbf{y})}c^{\pi}h_{\pi}\leq\sum_{(i,j)\in\mathcal{A}}\int_{0}^{x_{ij}^{\mathrm{f}}}t_{ij}(x)\,dx+\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}^{\mathrm{f}}}t_{l}(v)\,dv+\sum_{\pi\in\Pi(\mathbf{y})}c^{\pi}h_{\pi}^{\mathrm{f}}+\sum_{l\in\mathcal{C}}M_{l}(1-y_{l})$ (22)
Therefore, we have approximated cut (9j) in problem (9) with the combination of cuts (13), (17), and (22). The auxiliary variables $\eta_{ij}$ and $\psi_{l}$ are constrained by the OA cuts (13) and (17), which are iteratively added at each linearization point encountered during the algorithm. This approach transforms the nonlinear value function constraint into a collection of linear constraints that can be efficiently handled by MILP solvers. The final linearized HPR will be the following:
$\min_{\mathbf{y,x,v},\boldsymbol{\eta},\boldsymbol{\psi}}\quad$ $-\sum_{l\in\mathcal{C}}p\cdot v_{l}+w\cdot\sum_{(r,s)}\left(D_{rs}-\sum_{\pi\in\Pi_{rs}(\mathbf{y})}h_{\pi}\right)$ (23a)
@@
$0\leq y_{l}\leq 1,$ $\quad\forall l\in\mathcal{C}$ (23c)
@@
$v_{l}=\sum_{\pi\in\Pi(\mathbf{y})}\phi_{l}^{\pi}h_{\pi},$ $\quad\forall l\in\mathcal{C}$ (23g)
@@
$\int_{0}^{v_{l}^{1}}t_{l}(v)\,dv+t_{l}(v^{1})\cdot(v_{l}-v_{l}^{1})\leq\psi_{l},$ $\quad\forall\,\mathbf{v}^{1}$ (23k)
$\sum_{(i,j)\in\mathcal{A}}\eta_{ij}+\sum_{l\in\mathcal{C}}\psi_{l}+\sum_{\pi\in\Pi(\mathbf{y})}c^{\pi}h_{\pi}$
$\leq\sum_{(i,j)\in\mathcal{A}}\int_{0}^{x_{ij}^{\mathrm{f}}}t_{ij}(x)\,dx+\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}^{\mathrm{f}}}t_{l}(v)\,dv+\sum_{\pi\in\Pi(\mathbf{y})}c^{\pi}h_{\pi}^{\mathrm{f}}+\sum_{l\in\mathcal{C}}M_{l}(1-y_{l})$ $\quad\forall\,(\mathbf{x}^{\mathrm{f}},\mathbf{v}^{\mathrm{f}},\mathbf{h}^{\mathrm{f}})\in\mathcal{F}$ (23l)
@@ 5.2 Column Generation
Formulation (23) contains one path flow variable $h_{\pi}$ per feasible EV path $\pi\in\Pi(\mathbf{y})$. In realistic networks, $|\Pi(\mathbf{y})|$ is a large set. Following the approach used by [22], we apply the CG method to efficiently solve the linear programming (LP) relaxation. We consider a restricted set of feasible paths $\bar{\Pi}\subseteq\Pi(\mathbf{y})$, and denote the corresponding restricted path sets as $\bar{\Pi}_{rs}\subseteq\Pi_{rs}(\mathbf{y})$ for all OD pairs $(r,s)\in\mathcal{Z}^{2}$. Problem (23) can be written as follows:
$\min_{\mathbf{y},\mathbf{x},\mathbf{v},\mathbf{h},\boldsymbol{\eta},\boldsymbol{\psi}}\quad$ $-\sum_{l\in\mathcal{C}}p\cdot v_{l}+w\cdot\sum_{(r,s)}\left(D_{rs}-\sum_{\pi\in\bar{\Pi}_{rs}}h_{\pi}\right)$ (26a)
@@
$0\leq y_{l}\leq 1,$ $\quad\forall l\in\mathcal{C}$ (26c)
$v_{l}\leq M\cdot y_{l},$ $\quad\forall l\in\mathcal{C}$ (26d)
$x_{ij}=\sum_{\pi\in\bar{\Pi}}\delta_{ij}^{\pi}h_{\pi},$ $\quad\forall(i,j)\in\mathcal{A}$ (26e)
$\sum_{\pi\in\bar{\Pi}_{rs}}h_{\pi}=d_{rs}(\mathbf{y}),$ $\quad\forall(r,s)\in\mathcal{Z}^{2}$ (26f)
$v_{l}=\sum_{\pi\in\bar{\Pi}}\phi_{l}^{\pi}h_{\pi},$ $\quad\forall l\in\mathcal{C}$ (26g)
$\int_{0}^{x_{ij}^{1}}t_{ij}(x)\,dx+t_{ij}(x^{1})(x_{ij}-x_{ij}^{1})\leq\eta_{ij},$ $\quad\forall\mathbf{x}^{1}$ (26h)
$\int_{0}^{v_{l}^{1}}t_{l}(v)\,dv+t_{l}(v^{1})(v_{l}-v_{l}^{1})\leq\psi_{l},$ $\quad\forall\mathbf{v}^{1}$ (26i)
$\sum_{(i,j)\in\mathcal{A}}\eta_{ij}+\sum_{l\in\mathcal{C}}\psi_{l}+\sum_{\pi\in\bar{\Pi}}c^{\pi}h_{\pi}$
$\leq\sum_{(i,j)\in\mathcal{A}}\int_{0}^{x_{ij}^{\mathrm{f}}}t_{ij}(x)\,dx+\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}^{\mathrm{f}}}t_{l}(v)\,dv+\sum_{\pi\in\bar{\Pi}}c^{\pi}h_{\pi}^{\mathrm{f}}+\sum_{l\in\mathcal{C}}M_{l}(1-y_{l}),$ $\quad\forall(\mathbf{x}^{\mathrm{f}},\mathbf{v}^{\mathrm{f}},\mathbf{h}^{\mathrm{f}})\in\mathcal{F}$ (26j)
$h_{\pi}\geq 0,$ $\quad\forall\pi\in\bar{\Pi}$ (26k)
@@
Since the objective function aims to maximize path flows $h_{\pi}$ and charging flows $v_{l}$, we reformulate the corresponding constraints as inequalities to expose the sign of their associated dual variables. Specifically, the demand satisfaction constraint (26f) becomes $\sum_{\pi\in\bar{\Pi}_{rs}}h_{\pi}\leq d_{rs}(\mathbf{y})$, the link flow definition (26e) becomes $\sum_{\pi\in\bar{\Pi}}\delta_{ij}^{\pi}h_{\pi}\leq x_{ij}$, and the charging flow definition (26g) becomes $\sum_{\pi\in\bar{\Pi}}\phi_{l}^{\pi}h_{\pi}\geq v_{l}$. We derive dual variables associated with the constraints to determine the reduced cost of adding new paths to the LP. Denote by $\sigma_{rs}\geq 0$ the dual variables for the demand satisfaction constraints (26f), and by $\zeta_{ij}\geq 0$ the dual variables associated with link flow constraints (26e). Given an OD pair $(r,s)\in\mathcal{Z}^{2}$ and a path $\pi\in\Pi_{rs}^{\mathrm{f}}$, the reduced cost $\bar{c}_{\pi}$ for the path flow variable $h_{\pi}$ is computed as follows:
$\bar{c}_{\pi}=-\sigma_{rs}+\sum_{(i,j)\in\mathcal{A}}\delta_{ij}^{\pi}\zeta_{ij}-\sum_{l\in\mathcal{C}}\phi_{l}^{\pi}\gamma_{l}+c^{\pi},$ (27)
@@
where $\gamma_{l}\leq 0$ represents the dual variables for the charging station flow constraints (26g), and $c^{\pi}=\sum_{l\in\mathcal{C}}(c+\tau)\cdot\gamma_{l}^{\pi}$ is the path-dependent charging cost. Thus, to identify new paths that have negative reduced cost, one solves, for each OD pair $(r,s)$, a battery constrained shortest path problem with link costs given by dual variables $\zeta_{ij}$, station flow costs $\gamma_{l}$, and charging costs $(c+\tau)\cdot\gamma_{l}^{\pi}$. Since we have the expanded network we can solve the constrained shortest path problem (CSPP) efficiently. If paths with negative reduced costs are found, these are introduced into the restricted path set $\bar{\Pi}$ to refine the problem iteratively. This process continues until no paths with negative reduced cost are found. Algorithm 2 shows the CG steps to iteratively solve problem (26) using a limited set of paths and updates it by identifying and adding new paths with negative reduced cost.